gabriel could drive the 200 miles to the seashore in half the time it took martha to drive the 650 miles to the mountains. Find the speeds and times of both if martha's speed was 25 miles per hour greater than gabriel's speed.
Answer by mananth(16946)


gabriel could drive the 200 miles to the seashore in half the time it took martha to drive the 650 miles to the mountains. Find the speeds and times of both if martha's speed was 25 miles per hour greater than gabriel's speed.
Gabriel's distance 200 miles
Martha's distance = 650 miles
Let Gabriel's speed be x
Martha's speed = x+25
Martha's time = 650/(x+25)
Gabriel's time = 200/x
200/x = (1/2)*650/(x+25)
200*2/650 =x/(x+25)
13x=8x+200
5x = 200
x=40
Gabriel's speed 1s 40 mph
Time = 200/40 = 5 hours
Martha's speed = 65 mph
Time = 650/65 = 10 hours